Determine the mass of a sample of Vitamin D, C 27 H 44 O 2 , if it contains 2.4 x 10 22 hydrogen atoms.


1
Expert's answer
2021-05-24T07:18:30-0400

Solution:

One mole of any substance contains 6.022×1023 atoms/molecules.

Hence,

(2.4×1022 hydrogen atoms) × (1 mol / 6.022×1023 atoms) = 0.03985 mol H


C27H44O2 → 44H

According to the equation above: n(C27H44O2) = n(H)/44

n(C27H44O2) = (0.03985 mol) / 44 = 0.000906 mol

The molar mass of C27H44O2 is 400.64 g/mol.

Hence,

(0.000906 mol C27H44O2) × (400.64 g C27H44O2 / 1 mol C27H44O2) = 0.363 g C27H44O2


Answer: 0.363 g of a sample of Vitamin D (C27H44O2).